### v1

# Code Difference Guided Fuzzing for FPGA Logic Synthesis Compilers via Bayesian Optimization

Abstract
Field Programmable Gate Arrays (FPGAs) play a crucial role in Electronic Design Automation (EDA) applications, which have been widely used in safety-critical environments, including aerospace, chip manufacturing, and medical devices. A critical step in FPGA development is logic synthesis, which enables developers to translate their software designs into hardware net lists, which facilitates the physical implementation of the chip, detailed timing and power analysis, gate-level simulation, test vector generation, and optimization and consistency checking. However, bugs or incorrect implementations in FPGA logic synthesis compilers may lead to unexpected behaviors in target wapplications, posing security risks. Therefore, it is crucial to eliminate such bugs in FPGA logic synthesis compilers. The effectiveness of existing works is still limited by its simple, blind mutation strategy. To address this challenge, we propose a guided mutation strategy based on Bayesian optimization called LSC-Fuzz to detect bugs in FPGA logic synthesis compilers. Specifically, LSC-Fuzz consists of three components: the test-program generation component, the Bayesian diversity selection component, and the equivalent check component. By performing test-program generation and Bayesian diversity selection, LSC-Fuzz generates diverse and complex HDL code, thoroughly testing the FPGA logic synthesis compilers using equivalent check to detect bugs. Through three months, LSC-Fuzz has found 16 bugs, 12 of these has been confirmed by official technical support.
Index Terms:
FPGA, Logic Synthesis Compiler, Equivalent Check
I Introduction
In the realm of digital design and engineering, Hardware Design Language (HDL) logic synthesis tools are crucial, acting as a conduit from conceptual hardware models to executable digital circuits [1, 2, 3]. These tools allow engineers to convert high-level hardware functionality descriptions, articulated in HDL codes like VHDL or Verilog, into gate-level blueprints suitable for silicon chip fabrication. As shown as Figure 1, this conversion is essential for creating complex Integrated Circuits (ICs), encompassing CPUs, GPUs, and specialized ASICs.
Figure 1: The conversion from HDL code to netlist
Within the context of Intel’s FPGA Industrial Solutions Playbook, logic synthesis assumes a critical role, enabling these devices to seamlessly bridge and integrate with legacy industrial systems. This not only facilitates the modernization of existing infrastructure but also guarantees that systems can be promptly updated or reconfigured as needs evolve—eliminating the necessity for extensive hardware modifications. This adaptability is paramount in industries characterized by rapid technological advancement and where system downtime incurs significant costs [4, 5, 6]. However, logic synthesis tools are susceptible to bugs, stemming from the complexity of implementing extensive optimizations to fulfill stringent power consumption and timing constraints [7]. If unexpected behavior occurs in FPGA logic synthesis tools, may cause further finance loss and safety accident in legacy industrial systems. Thus it is important to keep the accuracy and safety of FPGA logic synthesis tools.
Recently, several methods have been developed to thoroughly test FPGA logic synthesis tools. Although they contribute significantly to testing FPGA logic synthesis tools, there remain some shortcomings. The optimal method for testing FPGA logic synthesis tools is LegoHDL [8], which generates a CPS model guided by a AST and uses HDLcoder to convert the CPS model to HDL code for FPGA logic synthesis testing. However, LegoHDL still has some disadvantage, such as low testing efficiency, and its random parameter configuration does not allow for the generation of more diverse test cases. Another testing method is Verismith [7], which uses an Abstract Syntax Tree (AST) generating method to create test cases and identify bugs. Due to the corpus’s shortcomings, Verismith can only generate less diverse and more repetitive codes. Despite its success in finding many bugs in FPGA logic synthesis compilers, Verismith’s effectiveness is still limited by its simple corpus.
We formulate our bug-finding process as an optimization problem whose goal is to maximize the difference between a variant and the seed program. By generating substantially diverse variants, we aim at testing more optimization strategies that may not be exercised otherwise. To address this issue, we propose a guided mutation strategy based on Bayesian optimization called LSC-Fuzz to detect bugs in FPGA logic synthesis compilers.
Specifically, LSC-Fuzz consists of three components: the test-program generation component, the Bayesian diversity selection component, and the equivalent check component. The test-program generation component, first collects coverage data in the seed CPS model. Then the insertion point is determined based on the data coverage collection information, and assertion statements are randomly inserted within the insertion point. Based on the coverage information of the seed model, LSC-Fuzz ensures that the inserted assertion statements are always true, and puts subsequent modules in the seed model into the true branch to ensure that they are still executed. Then LSC-Fuzz randomly select some modules to form a module sequence that conforms to the grammatical rules and add them to the newly created fake branch. The Bayesian diversity selection component first enriches the diversity of generating HDL code by adding specific timing relationship modules to the CPS model. The diversity of generated codes is then ensured by defining procedural distances and selecting LSC-Fuzz through Bayesian selection. By performing test-program generation and Bayesian diversity selection, LSC-Fuzz generates diverse and complex Verilog design file code. Finally, through the equivalent check component, if LSC-Fuzz detects inconsistencies in the output of the variant HDL code with the original HDL code, it may indicate the presence of bugs in the logic synthesis tools.
The experiments demonstrate that, within the same testing period, LSC-Fuzz significantly outperforms existing approaches, in the detection bugs of logic synthesis tools. Specifically, in three months, LSC-Fuzz has found 16 bugs, 12 of these has been confirmed by official technical support. In summary, the main contributions of this work are as follows:
  • We introduce LSC-Fuzz, a guided mutation strategy based on Bayesian optimization to uncover bugs in FPGA logic synthesis compilers. By performing test-program generation and Bayesian diversity selection, LSC-Fuzz generates diverse and complex HDL code, thoroughly testing the FPGA logic synthesis compilers using equivalent check to detect bugs.
  • Extensive experiments are conducted to evaluate the bug-finding capability of LSC-Fuzz. LSC-Fuzz has detected 16 valid bugs, of which 11 bugs have been confirmed as new bugs by official technical support.
  • We release LSC-Fuzz as a replication package for FPGA logic synthesis tool testing to foster reproducibility and facilitate further studies [9].
Figure 2: The process of logic synthesis
II Background and Motivation
II-A Preliminaries
Logic synthesis tools play a crucial role in FPGA development by converting high-level design descriptions into gate-level representations that are optimized for FPGA hardware. These tools allow designers to concentrate on the algorithmic and architectural elements of their projects, handling the transformation of abstract specifications into logic circuits that can be implemented on hardware. They enhance the design by optimizing for speed, area, and power consumption, thus shortening the time-to-market and boosting the efficiency of FPGA-based applications. Additionally, these tools facilitate iterative design processes by enabling developers to modify their designs based on feedback from simulations and real-world tests, thereby ensuring that the final product achieves the best possible performance and resource efficiency.
The process as shown as Figure 2, begins with the input of a design in a Hardware Description Language (HDL), such as Verilog or VHDL. These HDL scripts describe the circuit’s functionality but do not specify the physical implementation details. FPGA logic synthesis tools take these descriptions and perform optimizations and transformations, converting them into a mapped netlist of generic logic elements, such as lookup tables (LUTs) and flip-flops, which directly correspond to the resources available on the target FPGA.
Following the initial generation of the netlist, further optimizations are carried out to meet the specific requirements of the design. This includes optimizing the design for speed, power consumption, and area utilization. The FPGA logic synthesis tool applies various algorithms to rearrange, combine, or simplify logic to enhance performance and efficiency. Timing analysis is also performed to ensure that the circuit meets the desired operational frequency, and adjustments are made to the placement of logic elements and routing paths to minimize delays and avoid timing violations.
Finally, the logic synthesis process integrates the optimized netlist with the physical layout on the FPGA. This stage, known as place and route, involves placing the logic elements onto the physical grid of the FPGA and routing the interconnections between them. The output of this stage is a configuration file, often in the form of a bitstream, which can be loaded onto the FPGA to configure its logic blocks and interconnects to replicate the functionality defined by the HDL code. This file is crucial as it represents the final implementation of the design that will operate on the FPGA hardware.
II-B Illustrative Examples
We present a real confirmed bug that existing methods failed to detect. In this example, we demonstrate the effectiveness of our specific code generation approach. This bug involves a Vivado miscompilation issue, illustrated in Figure 3, which displays certain modules post-synthesis that trigger the error. Due to the complexity of the synthesized model diagram, which typically comprises thousands of modules, we focus solely on illustrating the root cause of the bug. The model generation process is outlined as follows:
Variable Profiling and Assertion Insertion: LSC-Fuzz profiles the variable valuations at all CPS points of the seed CPS model, followed by the insertion of an assertion statement. For instance, an If block is added where, based on the valuation of variables at specific CPS insertion points, the conditional predicate of this If block is enforced to always evaluate as true. Since the false condition of the If block is never executed, it introduces no side effects on the CPS model. This unused branch provides a new space to insert additional blocks, thus increasing the complexity of the generated code.
Subsystem Wrapping and HDL Conversion: Unlike procedural languages, both the true and false branches of the CPS model are encapsulated within a new subsystem. This subsystem, integral to the model, is then converted into HDL code by the HDL Coder. Throughout this process, all subsystems are translated into independent HDL code files and linked to the main file through references to ensure proper functional execution.
During synthesis of this example, the compiler crashed unexpectedly. Logically, the code within the else branch should not influence the circuit’s behavior, meaning that the generated code should be logically equivalent to the code prior to the insertion of the If block. However, this example triggered a compiler crash. The underlying issue was that the FPGA logic synthesis tool failed to appropriately manage the logic path containing the clock cycle reset signal during optimization. This failure caused the optimization algorithm to erroneously retain and execute the else branch code, which should have been disregarded. Consequently, under certain extreme timing conditions, the else branch was incorrectly selected, leading to an infinite loop and ultimately causing a crash.
Figure 3: Real FPGA logic synthesis compiler Bug found by LSC-Fuzz
III LSC-Fuzz Model
III-A Overview
Figure 4: The framework of LSC-Fuzz
The framework of LSC-Fuzz is illustrated in Figure 4 and Algorithm 1. LSC-Fuzz consists of three components: the test-program generation component, the Bayesian diversity selection component, and the equivalent check component. The basic idea of LSC-Fuzz is to generate diverse and equivalent HDL test case to exercise FPGA logic synthesis compiler thoroughly with equivalent test. Specifically, the test-program generation component (lines 10–15) profiles variable valuations of the CPS model and inserts assertion statements (e.g., If/While blocks) that always evaluate to be true at arbitrary points between two blocks in the seed CPS model. Hence, the outputs of the blocks before the assertion statements can be directly passed to their subsequent blocks in the true branches to preserve the equivalence of the CPS variants. LSC-Fuzz treat the body of unexecuted branches as a new space and fill it with randomly generated modules to generate new CPS models (Figure 4(a)). The Bayesian diversity selection component (lines 16–20) uses HDL Coder convert CPS model to HDL code and use bayes selector to effectively select test case to ensure the diversity of HDL code (Figure 4(b)). It generates complex HDL code with complicated timing relationship for triggering deeper bugs. Finally, LSC-Fuzz utilizes equivalent test (lines 1–9) to validate all HDL testcase (Figure 4(c)). If the output of a generated HDL code is different from the HDL code generated by seed CPS model, it is deemed to trigger a FPGA synthesis compiler bug.
Input: HDL Coder $L$, Seed CPS model $C$, CPS parameters $P$, FPGA logic Synthesis tool $S$, HDL Code $H$, Mutant number MAX-ITER
Output: Reported Bug
1
2
3Function TEST($S$, $C$, $P$, $L$, $H$):
4 $H\leftarrow L.\text{Compile}(C).\text{Execute}(P)$;
5 $O\leftarrow S.\text{Compile}(H)$;
6 for $i=1$ to MAX-ITER do
7
8    $H^{\prime}\leftarrow\textnormal{{TPC(}}\textnormal{\emph{C, P, L}}\textnormal{{)}}$;
9    $H^{\prime\prime}\leftarrow\textnormal{{BDS(}}\textnormal{\emph{H', S, H}}\textnormal{{)}}$;
10    $O^{\prime}\leftarrow S.\text{Compile}(H^{\prime\prime})$;
11    if $O^{\prime}\neq O$ then
12
13       return ReportBug(S, $H^{\prime\prime}$, P);
14
15
16
17Function TPC($C$, $P$, $L$):
18 $V\leftarrow\text{Profile}(C,P)$;
19 $a\leftarrow\text{GenAssertion}(C,P,V)$;
20 $C^{\prime}\leftarrow\text{FillSegment}(C,a)$;
21 $H^{\prime}\leftarrow L.\text{Compile}(C^{\prime}).\text{Execute}(P)$;
22 return $H^{\prime}$;
23
24Function BDS($H^{\prime}$, $S$, $H$):
25 $B\leftarrow\text{BayesSelector}(H^{\prime},H)$;
26 if $B\neq\text{FALSE}$ then
27
28    $H^{\prime\prime}\leftarrow\text{Optimize}(H^{\prime},S)$;
29 return Reported Bug $H^{\prime\prime}$;
Algorithm 1 LSC-Fuzz
III-B Test-Program Generation Component
The Test-Program Generation component (Function TPC from lines 10 to 15 in Algorithm1) systematically reshapes the control-flow structure at the state-machine level through two coordinated mechanisms: structural metamorphosis via path cloning and complexity amplification through dead-state injection.
Structural Metamorphosis via Path Cloning
Rather than merely inserting surface-level conditional branches, LSC-Fuzz fundamentally reconstructs the finite state machine (FSM) topology of the seed program by cloning feasible execution paths and embedding them within conditionally inhibited transitions. This metamorphosis preserves functional equivalence while exponentially expanding the state space that the logic synthesis compiler must navigate.
Specifically, LSC-Fuzz begins by profiling the reachable state space of the seed FSM model, capturing the invariant properties at each state transition point. Using Simulink’s Signal Range Coverage tool, we extract the state variable valuations across all feasible paths under the complete input domain. For each identified stable state $s_{i}$ in the original FSM, we synthesize a metamorphic transition structure that duplicates the state’s forward paths while maintaining deterministic execution through carefully crafted guard conditions.
Figure 5: State-space expansion through metamorphic transitions in logic synthesis
The metamorphic transformation adheres to structural constraints that ensure each cloned path remains functionally equivalent to its original. For any state variable $x$ participating in the transformation, we require that $\tau(x)\neq\text{boolean}\wedge\sigma(x)\neq\text{discrete}\wedge\psi(x)\text{ remains invariant within }T$, where $\tau(x)$ denotes the variable’s data type, $\sigma(x)$ its sampling characteristics, and $\psi(x)$ the state invariant that must hold throughout the sampling period $T$. This constraint formulation ensures that our metamorphic transitions operate on substantive state variables rather than simple boolean flags, maintain continuous or variable-rate sampling to preserve temporal dynamics, and guarantee that state invariants remain intact across all execution paths.
By constructing these metamorphic state transitions where the original execution path is embedded within a deterministically true branch, we ensure that the functional behavior remains unchanged while the structural complexity increases dramatically. This forces the logic synthesis compiler to reason about significantly more complex control-flow graphs without introducing any actual behavioral modifications.
Complexity Amplification through Dead-State Injection
Within the unreachable branches created by our metamorphic transitions, LSC-Fuzz strategically injects synthesized dead states that further amplify the structural complexity. These dead states, while never executed during normal operation, force the synthesis compiler to consider additional optimization paths and state encoding strategies.
The dead-state injection process follows rigorous constraints to maintain model validity. For each injected module $x$ within the dead-state set $\mathcal{M}$, we ensure that its data type $\tau(x)\in{\mathcal{F},\mathcal{I},\mathcal{C}}$ (representing float, integer, and complex types respectively) aligns with the surrounding state variables, preventing type inconsistencies that could trigger compiler errors. Furthermore, we require that the sampling time $\sigma(x)\in{\mathcal{T}_{c},\mathcal{T}_{v}}$ maintains either continuous or variable characteristics, preserving the temporal coherence of the state machine even in unreachable regions.
Most critically, we enforce that no injected module creates combinational loops within the dead states. Formally, for any module $x\in\mathcal{M}$ and any preceding module $y\in\mathcal{P}(x)$ in its dependency chain, we verify that $\neg\mathcal{L}(x,y)$, where $\mathcal{L}$ detects potential feedback paths. This constraint prevents the introduction of unstable configurations that could cause synthesis failures or infinite optimization loops.
Through this systematic approach to dead-state injection, we populate the unreachable state space with syntactically valid but semantically inactive structures. As illustrated in Figure 5, this expanded state space compels the FPGA logic synthesis compiler to explore a substantially broader range of optimization strategies, from state encoding and minimization to resource allocation and timing optimization. The compiler must process these complex state machines without knowing a priori which states are reachable, thereby exercising optimization paths that would remain untested with simpler input programs.
This state-machine-level metamorphosis represents a fundamental departure from traditional mutation testing approaches that operate at the expression or statement level. By reshaping the control-flow topology while preserving functional equivalence, LSC-Fuzz creates test programs that stress the synthesis compiler’s ability to handle complex state machines—a critical requirement for modern hardware designs where FSM optimization directly impacts circuit performance and resource utilization.
III-C Bayesian Diversity Selection Component
The Bayesian diversity selection component (as shown as Algorithm2) first convert the CPS model variant generated by the Test-Program Generation component into HDL Code through HDL Coder. Secondly, the Bayesian diversity selection component selects program variants with higher diversity through program distance and Bayesian selector. Through the Bayesian diversity selection component LSC-Fuzz, more diverse program variants are selected for testing, which enriches the diversity of the program and increases the ability of test cases to trigger deeper bugs.
1
Input: Seed CPS model $c$, HDL Coder $L$, List of program variants $V$, Prior probabilities
Output: Selected variants for testing
2
3
4Function FBDS($c$, $L$, $V$, $P(V)$):
5 $H\leftarrow L.\text{Compile}(c)$;
6
7 $d(V_{i})\leftarrow\text{FDist}(H,V_{i})$;
8
9 foreach $V_{i}\in V$ do
10    $d(V_{i})\leftarrow\text{FDist}(H,V_{i})$;
11
12    $P(V_{i}|D)\leftarrow\text{FProb}(V_{i},d(V_{i}),P(V))$;
13
14
15  Selected $\leftarrow$ Sort and select top variants based on $P(V_{i}|D)$;
16
17 return SelectedVariants;
18
19
20Function Dist($H$, $V_{i}$):
21 return $\sqrt{(v_{1H}-v_{1V_{i}})^{2}+(c_{1H}-c_{1V_{i}})^{2}+(s_{1H}-s_{1V_{i}})^{2}}$;
22
23
24Function Prob($V_{i}$, $d$, $P(V)$):
25 $\text{numerator}\leftarrow P(D|V_{i})\cdot P(V_{i})$;
26
27 $\text{denominator}\leftarrow\sum_{j=1}^{|V|}P(D|V_{j})\cdot P(V_{j})$;
28
29 return $\frac{\text{numerator}}{\text{denominator}}$;
30
31
Algorithm 2 Bayesian Diversity Selection
The Bayesian selector is a decision-making tool based on Bayesian decision theory and is used to make optimal choices under uncertainty. It uses prior knowledge and newly acquired data to update probability estimates to guide decision-making. The basic strategy of Bayesian selectors is to select variants with high temporal complexity and program complexity. LSC-Fuzz first define the program distance $d$ between two programs as follow.
$d(P_{1},P_{2})=\sqrt{(v_{1}-v_{2})^{2}+(c_{1}-c_{2})^{2}+(s_{1}-s_{2})^{2}}$ (1)
Formula (1) defines the Euclidean distance $d(P_{1},P_{2})$ between two programs $P_{1}$ and $P_{2}$ based on three metrics: the number of net variables $v$, the number of connections $c$, and the number of process structures $s$. Each metric contributes to the squared difference in their respective counts between the two programs, summed and then square-rooted to produce the distance. This measure quantifies the structural similarity between two programs, where a smaller distance indicates greater similarity in their composition and complexity.
Then LSC-Fuzz utilize Bayesian inference. Prior probability $P(V_{i})$ for each variant $V_{i}$ represents our initial belief about its complexity. The likelihood $P(D|V_{i})$, indicates the probability of observing specific complexity metrics $D$ given a variant $V_{i}$. The posterior probability $P(V_{i}|D)$, which updates our belief in the complexity of each variant based on observed data, is computed using formula (2):
$P(V_{i}|D)=\frac{P(D|V_{i})\cdot P(V_{i})}{\sum_{j=1}^{N}P(D|V_{j})\cdot P(V_{j})}$ (2)
Specifically, given a variant $V_{i}$, the choice to accept the variant $V_{i}$ depends on two aspects: the program distance $P(D|V_{i})$ and timing complexity ratio $P(V_{i})$ of the program from the original program, and the program distance $P(D|V_{j})$ and timing complexity ratio $P(V_{i})$ of the program to other programs.
Timing complexity in digital circuits refers to the measurement of the time requirements for signal transmission and processing, with a focus on the maximum time taken for a signal to travel from one logic unit to another along the longest path, known as the critical path. This maximum delay, constrained by the circuit’s clock frequency and synchronization operations, directly impacts the circuit’s performance and operational frequency. The timing complexity is essential in ensuring that all signal operations within the circuit are completed within one clock cycle to avoid timing errors and performance bottlenecks. The calculation formula of timing complexity is as shown in the formula (3)
$P(V_{i})=\max\left(\sum_{i=1}^{n}t_{i}\right)$ (3)
where $P(V_{i})$ represents the maximum delay across the circuit, essentially the delay along the critical path; $n$ is the number of components on that path; $t_{i}$ denotes the delay contributed by the $i$-th component. This calculation ensures that all components complete their operations before the arrival of the next clock signal, thereby maintaining circuit integrity and efficiency.
In this way, LSC-Fuzz calculate the acceptance probability $P(V_{i}|D)$ of all variant $V_{i}$ and form these probabilities into a one-dimensional matrix. This simplifies the selection of variants. If the test program setting requires multiple variants, LSC-Fuzz select those with the highest acceptance probability from the matrix. Through Bayesian selectors, LSC-Fuzz efficiently generates more complex HDL code for variant test programs.
III-D Equivalent Check Component
Since the netlist files generated by logic synthesis are usually not directly comparable, LSC-Fuzz have chosen two verification methods to verify the consistency of the netlist file, Satisfiability Modulo Theories(SMT) formalize verification and simulation verification.
III-D1 SMT formalize verification
Formal verification with SMT solvers entails several steps, ranging from defining logical constraints to solving them. LSC-Fuzz compared different netlists using the SMT solver ABC, provided by Yosys. The first step in SMT verification is to clarify the constraints.
$\mathcal{P}(\mathcal{N})=\{\phi_{1},\phi_{2},\ldots,\phi_{n}\}$ (4)
Fomula (4) introduces $\mathcal{P}$, which represents the set of properties $\phi_{i}$ associated with a netlist $\mathcal{N}$. These properties include functional characteristics and constraints that the netlist must satisfy.
After defining the constraints, LSC-Fuzz use the solver to check whether the constraints are satisfied, as described in Formula (5).
$\mathcal{C}(\mathcal{N}_{\text{orig}},\mathcal{N}_{\text{var}})=\bigwedge_{i=1}^{n}(\phi_{i}^{\text{orig}}\leftrightarrow\phi_{i}^{\text{var}})$ (5)
where $\mathcal{C}$ denotes the conjunction of bi-conditional constraints between corresponding properties of the original $\mathcal{N}_{\text{orig}}$ and variant $\mathcal{N}_{\text{var}}$ netlists, ensuring that each property $\phi_{i}$ is equivalently satisfied in both.
The SMT solver performs these checks to confirm that the netlist file is consistent before and after simulation. If there is an inconsistency, LSC-Fuzz will report the problem.
Figure 6: The output of FPGA logic synthesis compilers
It is worth noting that although the variant file generated by LSC-Fuzz is theoretically consistent with the output of the original file. However, they are not completely consistent in the SMT equivalence check, because the parts that will not be executed will also be reflected in the netlist file. Therefore, the SMT check has a false positive problem when verifying the variant file and the original file. To solve this problem, LSC-Fuzz chose to use simulation verification to further verify the FPGA logic synthesis compiler
III-D2 Simulation verification
This component consists of two sub-processes using different simulation tools and comparing original and variant model outputs. The output of the simulation tool is a waveform containing high and low levels(as shown in the Figure 6), so LSC-Fuzz print the output as a 0/1 level signal format and compare them bit by bit.
For different simulation tools, LSC-Fuzz use the same file to generate netlist files through different FPGA logic synthesis compilers, and use the same stimulus file to simulate them to compare their outputs. Fomula (6) represents the execution of simulation $\mathcal{S}_{j}$ on netlist $\mathcal{N}$ using stimuli $\sigma$, resulting in output $\mathcal{O}_{j}$. Each simulation tool might use a slightly different engine or algorithm, and $\mathcal{O}_{j}$ captures the output of the $j$-th tool.
$\mathcal{S}_{j}(\mathcal{N},\sigma)=\mathcal{O}_{j}\quad\text{for}\,j=1,2,\ldots,k$ (6)
After simulations, outputs are compared to assess the consistency of the netlist’s behavior across different simulation platforms based on Fomula (7). This formula ensures that all outputs $\mathcal{O}_{i}$ are identical for every pair of simulations, confirming that the netlist behaves consistently irrespective of the simulation tool used. If the output is inconsistent, LSC-Fuzz will report the issue. We will manually check the issues reported by LSC-Fuzz and report to developer.
$\text{Consistent}(\mathcal{O}_{1},\mathcal{O}_{2},\ldots,\mathcal{O}_{k})=\bigwedge_{i=1}^{k-1}\bigwedge_{j=i+1}^{k}(\mathcal{O}_{i}=\mathcal{O}_{j})$ (7)
For comparing original and variant model outputs, LSC-Fuzz prints the signal as a 0/1 electrical signal, and compares the simulation result output after the original CPS model is converted to HDL code and synthesized bit by bit with the simulation result output after the variant CPS model is converted into HDL code and synthesized. Theoretically the results should not change since the inserted module will not be executed and translated. But if the results are inconsistent, LSC-Fuzz will report the vulnerability. We will manually check these issues to find the root causes and submit them to the corresponding software developers.
IV Evaluation
In this section, four experiments are conducted to evaluate the effectiveness of LSC-Fuzz. Specifically, our evaluation aims at answering the following Research Questions (RQs).
RQ1: How is the bug-finding capability of LSC-Fuzz compared to the state-of-the-art methods?
RQ2: Can LSC-Fuzzdetect new FPGA logic synthesis compiler bugs?
RQ3: How effectiveness is the HDL code generation of LSC-Fuzz compared to the state-of-the-art methods?
RQ4: How effectiveness is the Bayes Selector of LSC-Fuzz?
In our experiments, RQ1 and RQ2 are used to evaluate the bug-finding capability of LSC-Fuzz compared to the state-of-the-art approaches. RQ3 and RQ4 are employed to evaluate the code generation strategy of the code generation component and the diverse variant generation component.
IV-A State of The Art Method
There are various testing methods available for FPGA logic synthesis compilers, such as VerilogHammer, Verismith, and LegoHDL. However, since VerilogHammer does not support the latest Verilog standard, we have selected Verismith and LegoHDL as our baseline methods.
IV-B Evaluation Setup
LSC-Fuzz has been developed using MATLAB and Python, and both our code and experimental data are publicly accessible on GitHub [9]. The evaluation of LSC-Fuzz was conducted on a computer running the Ubuntu 22.04 operating system, equipped with an Intel Core i9 CPU @ 2.10GHz, and 128GB of memory. Based on the recommendations of existing work, we focus on the size of generated HDL files being in the 700-100 line range [7].
In order to clearly illustrate the root cause of bugs, we utilize automated reduction approaches to simplify the HDL code that triggers the bug. This process enables developers to quickly comprehend and rectify bugs. Specifically, we employ a technique resembling the binary search method. Leveraging the AST extracted from the HDL code, we iteratively delete portions of the code until the erroneous use case can no longer be minimized. To avoid reporting duplicate bugs, we manually use failed assertions and back-trace to detect duplicates. When two bugs have same failed assertion or back-trace, we consider them as duplicates. Finally, we report the detected bugs which are verified as non-duplicates as new issues to developers.
IV-C Answer to RQ1
Approach. To evaluate the effectiveness of LSC-Fuzz, we compare the bug-finding capability of LSC-Fuzz with the state-of-the-art methods Verismith [7] and LegoHDL [8], since finding more bugs within a time period is the main objective of these approaches. In the experiment, we detect bugs on both the recently released version of FPGA logic synthesis compiler include Vivado, yosys, Iverilog and Quartus. We set a single testing period of two weeks for FPGA logic synthesis compiler, that is, every method tests these FPGA sythesis compilers for two weeks.
TABLE I: Bugs detected by LSC-Fuzz, LegoHDL, and Verismith
Known New Total
LegoHDL 2 1 3
Verismith 0 1 1
LSC-Fuzz 2 4 6
Results. As illustrated in Table I, bugs detected in our experiment are categorized into new bugs (New) and previously identified ones (Known), with the latter being duplicates from the bug repository. Notably, Table I demonstrates that LSC-Fuzz significantly outperforms Verismith and LegoHDL in bug detection capabilities. Over four weeks, LSC-Fuzz identified six bugs in these FPGA logic synthesis compilers, comprising two new bugs and four known ones. In contrast, LegoHDL detected three bugs—two known and one new—while Verismith identified only one known bug. The comparative analysis of bugs detected by Verismith, LegoHDL, and LSC-Fuzz in FPGA logic synthesis compilers is detailed in Figure 7. This occurs because LSC-Fuzz is capable of generating HDL test cases that are more complex than those produced by the baseline method. During the translation of the CPS model to HDL code by the HDL coder, each subsystem is converted into a separate HDL file with defined reference relationships. LSC-Fuzz enhances the complexity of these relationships by integrating complex code branches into the CPS model, thereby challenging the FPGA logic synthesis compiler to employ more sophisticated optimization strategies for the HDL files. Additionally, the Bayesian selector intensively tests the compiler by selecting increasingly complex HDL code scenarios for processing. All these lead to LSC-Fuzz’s bug discovery efficiency being better than the baseline method.
Conclusion. LSC-Fuzz significantly outperforms LegoHDL and Verismith for testing FPGA logic synthesis compiler. These results verify the bug-finding capability of LSC-Fuzz.
IV-D Answer to RQ2
Approach. Over a two-month period from May to July 2024, we extensively tested LSC-Fuzz to assess its bug-finding capabilities. Unfortunately, we could not validate the latest version of Quartus due to licensing issues. However, we employed the latest versions of other FPGA logic synthesis compilers for testing. Each identified issue with a logic synthesis compiler was reported to the corresponding software community, and we engaged with developers there to gather feedback.
Results. As depicted in Table II, our approach successfully identified 16 bugs within a three-month timeframe. The official websites of various logic synthesis tools, including Iverilog, Vivado, Yosys, and Quartus, have classified each issue as either Confirmed as New, Confirmed as Known, or Unconfirmed in Doubt (Pending Verification). New indicates bugs previously unrecognized by developers. Known refers to issues already identified as bugs by developers. Pending represents issues that developers believe can be circumvented through specific standardized procedures. To facilitate replication of our findings, we have made the HDL code files that triggered these bugs available on GitHub [9]. Notably, 4 new bugs were detected, which were not identified by previous methods. Figure 8 showcases the bug discovered by LSC-Fuzz. During the synthesis of HDL code, the compiler failed to correctly manage the reference relationships, leading to incorrect execution within the module positive. The Nonpositive block, correctly transformed into Verilog, reveals how intricate reference relationships can induce synthesis errors and potential compiler crashes. Existing methods like Verismith do not detect this vulnerability due to their limited code generation scopes that fail to produce complex reference relationships between blocks. This bug significantly impacts Vivado [10].
TABLE II: Vulnerability found by LSC-Fuzz
ID Summary Status Type Software
Yosys-1 support error for Aggregate Initialization and Replication Verified C Yosys
Yosys-2 Buffer Overflow Triggered by Inadequate Input Validation Verified C Yosys
Yosys-3 Incorrect State Machine Transition Due to Race Condition Verified C Yosys
Yosys-4 Memory Leak Arising from Unreleased Dynamic Allocations Pending M Yosys
Iverilog-1 SystemVerilog reference error Verified M Iverilog
Vivado-1 Synthesis crash under unsigned delay of specific symbol Verified C Vivado
Vivado-2 Synthesis faild and crush(HARTGLADDGen:regenerate(bool)) Verified C Vivado
Vivado-3 Vivado synthesis faild (ConstPropagte) Verified C Vivado
Vivado-4 Vivado synthesis failed (NANname) Verified C Vivado
Vivado-5 Synthesis failed caused by ConstProp::reconnect(NNet,NNet). Verified C Vivado
Vivado-6 Synthesis failed caused by inspection function. Verified C Vivado
Vivado-7 Reference relationship error casued synthesis failed Pending C Vivado
Vivado-8 Synthesis failed caused by NNet function Verified C Vivado
Vivado-9 Data Corruption from Improper Handling of Concurrency Verified M Vivado
Vivado-10 Inconsistent Data Integrity Caused by Faulty Synchronization Pending M Vivado
Quartus-1 Global RAM value reset inconsistencies Pending C Quartus
  • 1
    There are two types of status feedback from MathWorks on bug report (i.e., $Verified$ = newly confirmed bug, pendingpending = Pending verification). There are two types of bugs (i.e., Type) in our reported bugs: crash bugs ($C$) and miscompilation bugs ($M$).
Conclusion. LSC-Fuzz has proven effective in detecting bugs in FPGA logic synthesis compilers. Within just two months, we reported 16 valid bugs, four of which have been newly confirmed.
IV-E Answer to RQ3
Approach. We employ LSC-Fuzz, LegoHDL, and Verismith to generate HDL code test cases, each consisting of 700-1000 lines and totaling 1000 test cases. We compare their complexity and generation time using three key indicators derived from the Abstract Syntax Tree (AST) provided by Yosys. Our first metric, the count of blocks, involves tallying the blocks in files generated by both our model and the baselines, with a higher block count indicating greater complexity. The second metric, the count of lines or connections, suggests that a higher number of connections might increase the likelihood of encountering bugs. The third metric evaluates specific reference relationships, which when complex, are more likely to expose bugs.
Figure 7: Real logic synthesis compiler bug found by LSC-Fuzz, LegoHDL and Verismith
Figure 8: Real lofic synthesis compiler bug found by LSC-Fuzz
Results. Numbers of Blocks and Connections. Blocks and connections are vital components of HDL code, crucial for depicting its characteristics. As shown in Figure9, LSC-Fuzz generates a range of [40,49] blocks and [125,226] connections, surpassing the ranges produced by Verismith [7] ([25,37] blocks and [102,117] connections) and LegoHDL ([30,41] blocks and [120,210] connections). This increased complexity arises as LSC-Fuzz continuously selects more intricate HDL codes for testing, based on a Bayesian selector, enhancing its bug detection efficiency over the baseline methods. Specific Reference Relationships. In HDL design, complex reference relationships can challenge the compiler’s execution capabilities and optimize circuit design. To quantify this metric, we fixed the top-level design module and counted the number of modules it references. Figure9 shows that LSC-Fuzz’s range of module references exceeds those of Verismith [7] and LegoHDL. Specifically, LSC-Fuzz’s median reference count is 20% higher than LegoHDL’s and 33.3% higher than Verismith’s. On average, LSC-Fuzz’s reference count is 16.67% higher than LegoHDL’s and 25% higher than Verismith’s. The increase is attributed to LSC-Fuzz’s addition of more If branches, which are translated into independent files with reference relationships during translation, thereby complicating the application relationship. This complexity explains LSC-Fuzz’s ability to uncover issues that baseline methods miss, as it more thoroughly exercises the FPGA logic synthesis compiler with these complex relationships.
Conclusion. LSC-Fuzz outperforms baseline methods by generating more complex HDL test cases, enhancing its capability to detect bugs in FPGA logic synthesis compilers.
IV-F Answer to RQ4
Approach. To evaluate the impact of the Bayesian selector on HDL code generation, we conducted an ablation study comparing the complexity and generation time of HDL codes produced with and without the Bayesian selector. Each scenario involved generating 100 HDL codes, each ranging from 700-1000 lines.
Results. As shown as Figure 10, the generation of HDL code without the Bayesian selector proved to be 7% faster on average compared to scenarios where the selector was used. However, the HDL codes generated without the Bayesian selector contained approximately 18% fewer blocks and connections, indicating reduced complexity. Furthermore, these codes exhibited 10% fewer reference relationships, suggesting that even though the Bayesian selector does not directly alter HDL code, its absence noticeably decreases the intricacy of the generated codes. Thus, incorporating the Bayesian selector enhances the complexity of HDL test cases without significantly impacting generation speed.
Conclusion. The Bayesian selector effectively increases the complexity of generated test cases without compromising the efficiency of the generation process.
V Threats to Validity
Internal Threats. The primary threat to the internal validity of our method lies in the accuracy of LSC-Fuzz, as well as the reproduced models Verismith [7] and LegoHDL [8]. To mitigate errors in LSC-Fuzz, we meticulously double-checked our code. Additionally, we reproduced Verismith and LegoHDL by closely following the open-source code provided in their original publications, and we have thoroughly verified their accuracy. Furthermore, the model checking capabilities of Simulink HDL Coder can lead to increased processing times when generating large volumes of code. To address this, we optimized block placement to minimize model checking duration. Moving forward, we plan to intensify our testing of the FPGA logic Synthesis compiler by focusing on the generation of more complex test cases.
Figure 9: Blocks, connections and references generated by LSC-Fuzz, LegoHDL and Verismith
Figure 10: Blocks, connections and references comparison with Bayesian selector
External Threats. A significant threat to the external validity of LSC-Fuzz is the potential for triggering duplicate bugs during testing. To mitigate this, we have implemented strategies to reduce bug-triggering variants by excising blocks unrelated to the core issues, which enhances our ability to accurately understand and analyze bugs. Based on the developer’s verification, the bugs we reported are currently not duplicates, suggesting that our approach is effective in minimizing this threat.
VI Related Work
VI-A FPGA Logic Synthesis Testing
In the field of FPGA logic synthesis compiler testing, Verismith [7] and LegoHDL emerge as the primary methodologies for generating HDL code and testing logic synthesis tools. Verismith [7] functions as a Verilog program generator that crafts random behavioral Verilog code devoid of undefined values, based on predefined parameter configurations. Nevertheless, Verismith’s [7] capability to generate complex HDL code is limited as it exclusively produces Verilog, a single hardware design language.
To overcome these limitations and enhance the thoroughness of FPGA logic synthesis compiler testing, Xu et al. introduced LegoHDL [8], which aims to ensure the accuracy and correctness of FPGA logic synthesis processes. LegoHDL employs HDLcoder to generate HDL code by constructing a CPS model. It extracts the AST from the HDL code and orchestrates the construction of the CPS model to facilitate interaction between CPS model construction and HDL code generation. Finally, LegoHDL utilizes Metamorphic testing to synthesize the HDL code using various synthesizers and verify the equivalence of the generated netlist files. However, LegoHDL still faces challenges with insufficient diversity, and its singular CPS construction approach does not adequately address deeper bugs.
Another significant tool in this domain is VlogHammer [11], a Verilog fuzzer specifically designed for testing major commercial FPGA logic synthesis compilers and various simulators. As of now, VlogHammer [11] has identified approximately 75 defects. However, in contrast to Verismith [7], VlogHammer does not generate multi-model programs and lacks support for key behavioral-level Verilog constructs, such as always blocks. Similarly, VlogHammer [11] also limits its output exclusively to Verilog.
Additionally, random Verilog generators such as VERGEN [12] exist, producing behavioral-level Verilog by randomly assembling high-level logic blocks, including state machines, MUXes, and shift registers. These generators, however, depend on predefined definitions to construct their structures, leading to a limited diversity of generated configurations. This limitation reduces their ability to thoroughly test a wide spectrum of Verilog structure combinations.
In recent years, Large Language Models (LLMs) like the open-source CodeGen-16B, used in VeriGen [13] for test case generation, have shown promising results in FPGA logic synthesis compiler testing, boasting a 41% improvement in generating syntactically correct Verilog code over traditional pre-trained LLM methods. Despite this, VeriGen’s focus is not primarily on testing; its generation is limited to a dataset sourced from GitHub and textbooks, which may not thoroughly challenge FPGA compilers and lacks the capability to generate stimulus files, thereby restricting test accuracy. Other approaches, such as BetterV [14], are focused on optimizing Verilog code and reducing runtime. Currently constrained by the accuracy of the corpus and the scale of model generation, LLMs prove insufficient for effective HDL generation testing.
Furthermore, numerous efforts have been dedicated to enhancing the reliability of FPGA development compilers. For example, Yann Herklotz et al.[15] have concentrated on enhancing the stability of high-level synthesis (HLS) tools through formal verification, introducing Vericert, a formally verified HLS tool. Zewei Du et al.[16] have investigated the use of fuzz testing to comprehensively evaluate HLS tools by deploying a vast array of valid C programs. Despite these contributions, a gap remains in the quality and complexity of the generated files, predominantly focusing on a single programming language. Consequently, we have embarked on more comprehensive testing of FPGA logic synthesis compilers to affirm their stability and reliability.
VI-B Metamorphic testing in Compiler testing
Metamorphic testing generates variants of test cases considered equivalent, based on program inputs. It identifies bugs by comparing the outcomes of these variants [17, 18, 19, 20]. Metamorphic testing has been extensively verified with widely used compilation tools, including GCC, and this method has identified over 1,000 bugs [21, 22, 23, 24]. In compiler testing, three EMI-based Metamorphic testing mutation approaches are employed: Orion [21], Athena [22], and Hermes [23]. Orion specifically targets dynamically dead program regions, randomly pruning unexecuted statements to generate variant programs [21]. Athena inserts or deletes code in dead code regions under varying inputs [22]. Unlike Orion and Athena, which are limited to mutating dead code regions, Hermes [23] can alter both live and dead code regions to generate equivalent variants.
Metamorphic testing has also been applied in other fields. For example, SLEMI [25] proposed by Shaful et al. uses strategies such as deleting dead code to generate CPS variants to thoroughly test the Simulink compiler. Guo et al. proposed COMBAT [26], which uses the idea of live code mutation to generate CPS variants in order to more thoroughly test the Simulink compiler. In addition, the idea of Metamorphic testing has also been applied at the project level. Wu et al. proposed Mantra [27] to apply mutation testing to HDL and achieved good results.
VII Conclusions and Future Work
In this paper, we proposed a novel method for FPGA logic synthesis compiler testing, named LSC-Fuzz. This method incorporates a Test-Program Generation component and a Bayesian selection component, enabling the generation of diverse HDL code for testing the FPGA logic synthesis compiler. Over a three-month period, our method successfully identified 16 bugs, 12 of which have been confirmed by developers. Going forward, we plan to explore how to integrate large models to guide the generation of even more diverse test cases, further enhancing our testing of FPGA logic synthesis compilers.
Acknowledgment
This work was supported by the National Natural Science Foundation of China (No.62472062, No.62202079), the Dalian Excellent Young Project (2022RY35), the Fundamental Research Funds for the Central Universities (No.3132024257).